OpenStudy (anonymous):

2(x − 4) − 3 = −3(x − 5)

OpenStudy (anonymous):

2x-8-3+3x+15=0 5x+4=0 5x=-4 x=-4/5

OpenStudy (anonymous):

THANK YOU! I kept making the 3x negative

OpenStudy (anonymous):

can you click on the best response please :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Thank you